### v1

# StructLM: Towards Building Generalist Models for Structured Knowledge Grounding

Abstract
Structured data sources, such as tables, graphs, and databases, are ubiquitous knowledge sources. Despite the demonstrated capabilities of large language models (LLMs) on plain text, their proficiency in interpreting and utilizing structured data remains limited. Our investigation reveals a notable deficiency in LLMs’ ability to process structured data, e.g., ChatGPT lags behind state-of-the-art (SoTA) model by an average of 35%. To augment the Structured Knowledge Grounding (SKG) capabilities in LLMs, we have developed a comprehensive instruction tuning dataset comprising 1.1 million examples. Utilizing this dataset, we train a series of models, referred to as StructLM, based on the Code-LLaMA architecture, ranging from 7B to 34B parameters. Our StructLM series surpasses task-specific models on 14 out of 18 evaluated datasets and establishes new SoTA achievements on 7 SKG tasks. Furthermore, StructLM demonstrates exceptional generalization across 6 novel SKG tasks. Contrary to expectations, we observe that scaling model size offers marginal benefits, with StructLM-34B showing only slight improvements over StructLM-7B. This suggests that structured knowledge grounding is still a challenging task and requires more innovative design to push to a new level.
1 Introduction
Traditionally, users need to write programs to interface with structured data like tables, databases, knowledge graphs, etc. It requires the users to master the domain-specific language like SQL, SPARQL, etc. Recently, researchers have explored the possibility of automating the interface with natural language to enable potential use cases in question-answering Pasupat and Liang (2015); Zhong et al. (2017); Nan et al. (2022), summarization Parikh et al. (2020); Nan et al. (2021); Bao et al. (2018), and fact verification Aly et al. (2021); Chen et al. (2019); Gupta et al. (2020a), among others, all grounded to a structured knowledge source. This effort can lower the barrier for end users to access massive amount of structured data.
Figure 1: StructLM can ground on structured and unstructured knowledge to respond to human queries. The previous SoTA was attained by many different task-specific models like TAPEX Liu et al. (2021), USKG Xie et al. (2022), TableLLaMA Zhang et al. (2023), BINDER-Codex Cheng et al. (2022), etc. StructLM (a single model) beats the previous SoTAs on seven out of eighteen SKG tasks.
The previous work Yu et al. (2020); Liu et al. (2021); Xie et al. (2022); Zhang et al. (2023) has been mostly focused on building task-specific models for different tasks with rather limited generalization ability. Building a generalist structure knowledge grounding (SKG) system across a wide range of tasks proves to be challenging. This is mainly due to the heterogeneity of data format and use cases. We evaluated GPT-3.5-Turbo Jiang et al. (2023) on 18 SKG tasks and observed that its performance is on average 35% lower than the SoTA specialized models. It shows that the LLM’s ability on SKG is heavily overlooked during the pre-training phase.
Figure 2: Overview of StructLM. This figure illustrates the prompting structure of StructLM, highlighting its capability to process various forms of structured data beyond linearized data tables, including linearized database schemas and knowledge graphs. StructLM is also assessed on held-out tasks that bear similarity to groups of held-in tasks, but also differences that must be overcome.
In this paper, we explore the possibility of building a generalist model based on LLMs that can ground on diverse types of structure and unstructured knowledge to interface with humans. Specifically, we construct a large data set of over a million instruction-following examples, a majority of which is SKG data, along with additional general instruction-following data, which we find improves generalizability. We fine-tune models at three scales: 7B, 13B, and 34B, based on the Code-LLaMA family of code foundation models. When compared to USKG, we find that our 7B model surpasses these single-task models from on $11$ of $18$ tasks with our 34B model outperforming on $14$ of $18$. As depicted in Figure 1, StructLM achieves SoTA on $7$ out of 18 evaluated tasks, beating ChatGPT by a huge margin.
We study the performance of StructLM, namely whether the model experiences cross-task generalization benefits from the dataset mixture, and find that our multi-task model performs significantly better overall than single-task models of the exact same parameter scale. We also study the effect of different pretraining data on our finetuned performance to determine whether special pretraining regimes, such as code or math, contribute to effective SKG reasoning ability. We find that code pretraining is the most effective. We perform additional ablations to confirm our results and support our claims. Our contributions are:
  • We construct a large SKG instruction-tuning dataset with $1.1$ million samples. We train and release our 3 models that outperform the previous 3B USKG fine-tuned on individual tasks on a total of $14$ of $18$ tasks. StructLM also achieves SoTA results on $7$ of them.
  • We show that StructLM is able to show strong zero-shot generalization capability on unseen structure knowledge grounding tasks, which was not shown by previous models.
  • We find that scaling general instruction-tuning data improves generalization ability, and that code-pretrained base models indeed improve model performance on the SKG tasks.
2 Related Work
2.1 Solving SKG tasks
Structured knowledge, such as web tables, knowledge graphs, and databases, have long been the subject of study in knowledge grounding. However, SKG tasks have heterogeneous data formats which have inspired methods that leverage specific training setups to learn those representations. For example, PTab Liu et al. (2022) and MultiHiertt Zhao et al. (2022) learn the contextual representation of tabular data by incorporating semantic information through specific training methods or reasoning approaches. RASAT Qi et al. (2022a) integrates relation-aware self-attention with the Transformer seq2seq architecture and utilizes various relational structures to address SQL problems. TAPEX Liu et al. (2021) conducts pretraining over tabular/database data with the help of an SQL executor to provide supervision.
More recently, methods have begun to move away from these auxiliary task-specific structures. USKG Xie et al. (2022) were the first to unify many SKG tasks into a sequence-to-sequence format, allowing them to be aggregated into the same data mixture. StructGPT Jiang et al. (2023) represents a line of work that uses prompting frameworks on powerful LLMs to solve tasks with more robustness and accuracy. In contrast, our work examines open models and tries to assess their fundamental capabilities. TableLlama Zhang et al. (2023) is similar to our work but focuses only on tabular data, which we consider a subset of SKG.
USKG showed the benefit of the sequence-to-sequence format unification in more powerful language models, however, it was not able to show strong advantages to constructing a multi-task mix of SKG data over single-task models or task-specific training methods such as prefix-tuning in terms of performance. We seek to address this gap with our construction of an instruction-tuning SKG dataset with a large subset of USKG’s tasks, and the resulting model, StructLM, which does show strong multi-task performance.
2.2 LLMs with Instruction Tuning
Instruction-tuning (IT) has been popularized as a method to address the gap between training objectives and user goals in LLMs. This technique involves additional training of LLMs using pairs of instructions and outputs. IT enhances both the controllability and the predictability of the models, aligning them more closely with user expectations. Furthermore, recent studies such as FLAN Wei et al. (2022), UL2 Tay et al. (2023a), and LLAMA2 Touvron et al. (2023) have shown that IT can improve the performance of downstream tasks through multi-task learning across diverse data types. While FLAN-UL2 trains on a subset of 11 tasks from USKG, it also trains on many more unrelated language tasks. In our work, by focusing on SKG data, we hope to provide a focused study that can act as a reference for future work to improve performance on this task type.
2.3 Reasoning Capability in LLMs
Reasoning stands as a pivotal skill for LLMs in the development of real-world AI applications which would enable the autonomous completion of many thought-intensive tasks viewed traditionally to require human thinking, like programming or mathematical problem-solving Li et al. (2022). Recent studies Li et al. (2022, 2023c); Rozière et al. (2023); Azerbayev et al. (2023a) indicate that LLMs trained on code and mathematical datasets exhibit profound reasoning skills, and can even achieve performance on par with human levels. For example, Code-Llama Rozière et al. (2023), a foundation model trained on more programming data, has significantly improved reasoning capabilities across a variety of programming and mathematical benchmarks. Furthermore, Llemma Azerbayev et al. (2023a) continues to pretrain the Code-Llama model on a mix of scientific papers, math-related web data, and mathematical code. Its results show excellent reasoning capabilities on the MATH benchmark Hendrycks et al. (2021) and the ability to prove theorems without further fine-tuning. On the fine-tuning side, WizardMath Luo et al. (2023a), and WizardCoder Luo et al. (2023c) have shown the effectiveness of instruction tuning on reasoning capabilities, given high quality data.
In this work, we view structured data as a third testbed for a different kind of reasoning within LLMs. We posit that in addition to mathematical or logical reasoning, the ability to recognize and make use of patterns within a structured input indicates that a model has robust representations of relationships in data. These representations may serve as a strong prior for further reasoning downstream.
Figure 3: Breakdown of Structured Knowledge Types and Tasks. On the inside, we see a coarse breakdown of the different categories of structured inputs in our dataset. On the outside, we see the tasks that are represented for those structured knowledge types. We also have a considerable proportion of general instruction following data from SlimOrca Lian et al. (2023).
3 Method
3.1 Dataset Curation
Motivated by the goal of training a language model (LM) generally capable of a wide range of structured data tasks, we select a total of $25$ SKG tasks to study. We report results on $18$ held-in and $6$ held-out tasks, where each held-out task meant to roughly evaluate the generalization capability of a held-in task group. In total, our held-in training dataset contains approximately $700$k SKG examples. We describe the held-in dataset groups below.
Data to Text Generation. This group of datasets deals with the summarization or interpretation of structured data from tables to knowledge triples to formal languages. Their inclusion is motivated by the idea that useful LMs should be able to make sense of a wide variety of structured information and map it to meaning in natural language. The corresponding held-out dataset for this task group is intended to be WikiTableText.
Table based Question Answering. This group of datasets deals specifically with tabular data, optionally combined with text passages. LMs which are able to accurately answer questions and retrieve information from tables can be widely useful as assistants. The corresponding held-out dataset for this task group is SQA.
Knowledge-grounded Conversations. This group of tasks evaluates knowledge grounding in-conversation. Humans naturally interface with LMs is through chat, and enabling this capability can lower the barrier to accessing the information in stored structured data. These tasks track user intention through provided dialogue and ask the model to provide an answer to the latest question. The held-out dataset for this task group is CoSQL.
Fact verification. One common use case for tables is to reference facts. In addition to question answering, the ability to reliably determine if data in a table supports a statement signals the existence of a robust representation of the table’s data. The held-out dataset for this task group is InfoTabs.
SQL or domain-specific languages SQL is the language most commonly used to interface with structured data today. Understanding how to write SQL also requires understanding of abstractions of tables and how they are linked together. In other domain-specific languages, the MTOP task measures a model’s ability to parse a specification and generate an API call, which sees potential in LLM tool use (e.g., Qin et al. (2023)). The corresponding held-out dataset for this task group is intended to be BIRD Li et al. (2023b), which further tests SQL generation abilities.
Mathematical reasoning. An analysis of tabular data may also require performing quick mathematical computations over their contents. Performance on these datasets tells us how well models can combine both structured knowledge and mathematical reasoning. As there are currently a limited number of datasets that combine mathematical reasoning with SKG, this category includes just TabMWP in the held-in corpus. We set FinQA as a challenging held-out dataset analogue. Not only does it require financial domain knowledge, but it combines tabular information with long text passages, and requires the generation of mathematical code.
General instruction data. In addition to the SKG datasets within the held-in dataset mixture, we also included general instruction tuning data without any structured knowledge component, to maintain the instruction-following ability of our model. We use SlimOrca Lian et al. (2023), which is constructed from cleaned GPT-4 responses to a number of prompts from existing general large-scale instruction-following datasets. We detect no signs of data contamination for our held-out datasets based on our ablation results. We give a detailed overview of all dataset statistics in  Table 1.
3.2 Instruction Finetuning Approach
To instruction tune our model, each example in our dataset consists of a system prompt, instruction, input, and output. For all SKG data examples, we use the same system prompt. For each dataset, we write 10 instruction variations, which are randomized when constructing the training samples. For SKG data, the input is composed of a combination of a structured knowledge input and accompanying text that could be a question, statement, or anything that would be required to specify the task. The exact prompt format is provided in Figure 6.
3.3 Training and Evaluation Details
The base models for StructLM are the CodeLlama-Instruct family of models Rozière et al. (2023). We finetune all models with a batch size of $512$ for $3$ epochs on A800 gpus. This training setup is largely in line with community conventions, such as the settings used for the WizardLM Xu et al. (2023), WizardMath Luo et al. (2023a), and WizardCoder Luo et al. (2023c) models.
Overall Length Train Test
Dataset Input (avg) Output (avg) Count Input (max) Output (max) # trunc. Count Input (max) Output (max) # Trunc.
TabMWP 207.8 4.5 23059 709 33 0 7686 703 31 0
ToTTo 251.8 31.0 120761 2040 155 467 7700 2048 119 31
GrailQA 281.0 44.1 44337 884 134 0 6463 546 123 0
SQL2Text 122.3 18.1 5600 337 61 0 1034 245 38 0
MMQA 656.2 7.7 15688 2047 146 234 1501 2048 94 11
Spider 266.6 36.0 7000 1369 226 0 1034 453 146 0
KVRet 573.4 17.1 6288 1217 161 0 807 1147 82 0
HybridQA 700.4 6.8 62682 2047 91 200 3466 2048 79 6
SParC 276.3 32.6 12059 1417 226 0 1625 467 146 0
CompWebQ 1350.3 11.9 27639 2047 321 321 2816 2048 256 8
TabFact 660.1 4.6 92283 2045 5 2 12779 1687 4 0
WikiTQ 831.8 5.8 11321 2028 273 0 4344 2048 148 10
WikiSQL 689.2 7.1 56355 2047 518 16 15878 2048 244 1
FeTaQA 653.2 38.8 7326 1853 158 0 2003 1548 114 0
FEVEROUS 799.3 3.4 40669 2047 5 2052 4285 2048 4 195
MultiWOZ 777.2 154.5 56668 1656 196 0 7368 1344 185 0
DART 133.7 30.3 62659 406 258 0 5097 261 109 0
Logic2Text 166.1 26.9 8566 358 67 0 1092 347 60 0
MTOP 961.0 34.4 15667 1002 215 0 4386 990 113 0
SlimOrca 278.9 152.4 512069 2047 1808 0 - - - -
BIRD 439.8 63.3 9428 1992 347 99 1534 1214 386 0
CoSQL 287.4 34.9 9502 1640 226 0 1300 535 190 0
SQA 656.9 34.9 12275 1812 1012 2 3011 1725 769 0
Infotabs 276.9 3.7 16538 1009 5 0 5400 1105 4 0
WikiTableText 149.6 27.4 10000 313 97 0 2000 226 89 0
Finqa 1230.3 21.0 6251 2040 72 186 1147 2048 61 25
Table 1: Token sequence length statistics for each dataset in our train and test sets. Input and output statistics are in tokens. We report the number of examples which have been truncated in each dataset.
We follow the structured data linearization conventions in USKG Xie et al. (2022). However, we use a different truncation scheme as described below. During training, we maintain a maximum sequence length of $2048$. USKG notes that training and running inference with longer context is beneficial. As such, when truncating, we consider the combined token length of the prompt input and output label. We truncate only the structured knowledge portion of the input so that the example becomes at most $2048$ tokens long. As shown in the dataset statistics in Table 1, setting the max token length of the examples in our dataset to $2048$ allows nearly all examples to fit within the context window with rare truncations. We discard examples for which even this structured input truncation is insufficient (e.g. the output is too long). During inference, we set the input token length to $2048$, to allow even more structured information to be placed within the input context. We set the maximum generation length to 1024, which is sufficient for all correct responses in all datasets. For each model, including our single-task finetuned models, we choose the best performing checkpoint of the $3$ checkpoints produced at the end of each epoch.
4 Experiments
Dataset Metric SoTA ChatGPT Base ST UL2 TabLLaMA USKG StructLM (Ours) $\Delta$
Size - - - 7B 7B×$\times$18 20B 7B 3B×$\times$18 7B 13B 34B -
Held In
ToTTo BLEU 49.9 20.7 17.5 48.8 - 20.7 49.0 49.4 49.3 50.2 +1.2
GrailQA EM 77.1 9.3 1.0 77.0 - - 70.1 80.4 79.2 82.2 +12.1
SQL2Text Blec 94.8 88.6 82.9 95.2 - - 94.8 93.8 88.5 92.6 -1.0
MMQA F1 85.3 59.6 30.7 81.5 - - 85.3 85.2 86.0 88.1 +2.8
Spider EM 80.5 43.8 5.2 67.3 - - 71.8 72.4 74.1 74.6 +2.8
KVRet All Micro 67.9 52.9 39.5 70.9 - 48.7 67.9 72.6 69.5 69.3 +4.5
HybridQA Acc 68.4 23.7 2.3 58.4 61.0 27.6 59.4 59.2 59.1 61.1 +1.7
SParC EM 68.2 32.2 3.2 62.3 - - 61.5 61.9 64.9 63.4 +3.4
CompWebQ Acc 76.8 48.9 3.1 75.6 75.9 - 73.3 78.3 80.4 81.9 +8.6
TabFact Acc 93.0 62.4 0.0 79.6 87.1 82.5 83.7 80.8 84.7 86.6 +2.9
WikiTQ All Ex 65.9 24.8 0.2 45.7 54.6 31.6 49.3 50.1 53.4 55.7 +6.4
WikiSQL All Ex 93.0 31.5 0.4 86.5 87.3 41.6 86.0 88.7 87.2 87.6 +1.7
FeTaQA BLEU 39.0 7.4 5.6 33.8 35.8 39.0 33.4 36.0 35.6 37.5 +4.1
FEVEROUS Acc 85.6 57.8 58.4 78.1 85.6 72.3 82.4 84.4 85.0 85.7 +3.3
MultiWOZ Joint Acc 60.6 8.9 0.0 53.0 - - 55.4 54.5 53.0 53.8 -0.9
DART BLEU 52.0 59.0 54.6 60.3 50.4 - 46.7 62.2 61.4 61.8 +15.5
Logic2Text Blec 95.3 78.5 59.1 89.5 - - 91.4 88.9 90.1 89.1 -1.3
MTOP EM 87.5 1.4 0.0 77.4 87.5 - 86.8 81.2 81.6 82.1 -4.7
Average 74.9 39.5 20.2 68.2 - - 69.3 71.1 71.3 72.6 +3.3
Held Out
BIRD Acc 36.6 21.8 0.0 24.4 - 4.2 $\times$ 22.3 22.8 24.7 -
CoSQL EM 58.3 33.7 0.2 52.4 - 5.4 $\times$ 49.8 52.2 55.0 -
SQA Acc 70.5 18.7 2.3 60.4 70.1 2.0 $\times$ 49.7 36.1 44.2 -
Infotabs Acc 75.6 46.9 40.2 68.7 - 35.5 $\times$ 55.3 58.1 61.8 -
WikiTableText BLEU 33.7 3.8 5.7 39.8 - 10.2 $\times$ 8.3 9.3 8.8 -
Finqa Acc 71.1 31.4 1.7 79.7 - 18.6 $\times$ 27.3 25.6 36.2 -
Average 57.6 26.1 8.4 54.2 - 12.6 $\times$ 35.5 34.0 38.4 -
Table 2: The held-in evaluation results of our model against other baselines. TabLLaMA refers to the table fine-tuned LLaMA Zhang et al. (2023), the Held-out results are obtained by us. BASE refers to the 1-shot performance of CodeLlama-Instruct-7B. Single FT refers to the single-task finetuning result of CodeLlama-Instruct-7B on each dataset. XYZ denotes the state-of-the-art score achieved by us. All StructLM held-out results are 0-shot. Specifications as to how SoTA results are selected are given in Table 4.
Baselines
Firstly, to illustrate the current performance of language models on SKG tasks, we evaluate ChatGPT (GPT-3.5-turbo) and the base model CodeLlama-7B-Instruct under a 1-shot setting. Our prompting scheme, using the same linearized knowledge structures as in our held-in training, sees them struggle across the board with many tasks due to the unseen structure knowledge format. Although ChatGPT is superior on text-based tasks, its performance is lackluster on SKG tasks. Its gap with SoTA models is as significant as 35%.
Held-in Results
To evaluate the benefits of our instruction-tuning dataset mix, we also run single-task baseline (each a 7B model) on each task and report their individual performance. We again use CodeLLama-7B-Instruct as the base model for each, and match each single task model on the same number of epochs (3) that was used to train the multitask models, ensuring that each model has seen the same data the same number of times. We observe that our multi-task models outperform these single-task models on nearly every task, with some by a considerable margin of up to $7\%$. This demonstrates the effectiveness of our instruction tuning dataset and supports the presence of cross-task generalization within our dataset mix.
When compared to the single-task 18 USKG models, StructLM-7B can surpass USKG by a average of 2%. From a parameter-count perspective, each of the USKG models is a T5-3B model, which means over the entire held-in set, these results require 54B parameters. Our 7B StructLM in comparison can be viewed as being nearly 8x as parameter efficient while still surpassing USKG models on 11 of 18 datasets. It is worth noting that although our single-task (ST) models are more than double the size in parameters compared to USKG, they do not perform much better on average. This fact indicates that there may be significant unused model capacity that can be better utilized via more effective training regimes, such as our instruction tuning. Regarding FLAN-UL2-20B Tay et al. (2023b), which was also extensively trained on structure knowledge grounding tasks, StructLM outperforms it on 7 of the 9 mutually held-in datasets. Regarding TableLLaMA Zhang et al. (2023), which is a table-specialized LLM, StructLM can outperform it significantly across the board. Unlike StructLM, TableLLaMA cannot deal with structure beyond tables, which make it less generalizable than StructLM.
Held-out Results
On held out tasks, StructLM shows strong generalization performance, outperforming ChatGPT on 5 of 6 tasks. We note that our zero-shot prompting method for the FinQA Chen et al. (2021) dataset requires the model to generate custom code and is focused only on the financial domain, and the input to Infotabs Gupta et al. (2020a) is a different type of table structure unseen in the input data, with different labels for the boolean output categories than the choices from the held-in fact verification tasks. This generalization capability is non-existent in USKG models as the name of each held-in task is specified in their training inputs. When compared to TableLLaMA, StructLM also generalizes much better.
5 Ablation Studies
Figure 4: Effect of different pretraining curricula on SKG finetuning performance in relevant task groupings.
Purpose Train Eval FT Result
Schema task transfer Spider, SParC, Logic2Text Logic2Text 89.47 89.93
KT task transfer CompWebQ, WebQSP, GrailQa, Dart Dart 60.28 60.34
Table task transfer FetaQA, HybridQA, WikiTQ, TabMWP, ToTTo, MMQA, WikiSQL, KVRet, Tab Fact, Feverous, Infotabs TabFact, Feverous Infotabs 75.46 80.81
Summ. data type transfer ToTTo, Dart Dart 60.28 61.42
QA data type transfer CompWebQ, WikiSQL WikiSQL 85.49 86.36
Table 3: Cross task and cross datatype transfer results. FT is an average of single-task performance over the datasets in the Eval column.
Figure 5: Effect of general instruction-following data on held-out SKG dataset performance. Performance is measured as the average over evaluation metrics across all tasks within held-in or held-out groups. Note that the held-in performance experiences a milder dip compared to the held-out performance gains.
Effect of base model pretraining data. We ablate our choice of base model, Codellama-7b-Instruct, by finetuning the unspecialized Llama2-7b base model and Llemma, which is further pretrained on mathematical texts Azerbayev et al. (2023b). Intuitively, one might guess that coding ability has the most transferability to performance on the types of SKG tasks we are studying due to the symbolic nature of programming languages and code-writing scenarios. However, it is possible that other types of pretraining to boost reasoning ability, such as math, have even greater transferability.
Our ablation results in Table 6 can be broken down into groupings of tasks, as in Figure 4. Models pretrained on code indeed perform slightly better, and these gains are not necessarily limited to tasks which explicitly involve a grammatically regular input, or require the generation of code. Math pretraining does seem to improve the performance of the Llama2 base model, but not by as much as code pretraining. Overall, it seems that code pretraining is a useful step in training a performant model in this SKG setting, which may be due to conceptual similarity on certain tasks.
Effect of general instruction data in mixture We ablate our inclusion of approximately $500$k general instruction-following examples into the training data mixture to observe the importance of including this data. As we see in Figure 5, the held-in performance is relatively unaffected by the added general examples, but held-out performance improves significantly with more general data. Empirically, we also observe that when training a large volume of task-specific input and output formats, the model becomes less capable of following instructions on new tasks in a zero-shot setting. We hypothesize that training on this general mixture helps zero-shot performance because it can reduce overfitting to the task formats in the training set.
Cross-task and cross-format transferability We ablate the transferability of performance between input structure knowledge types and between output task types. To test this, we train a number of tasks together and compare them to their single-task finetuned models. Our results are indicative that there is cross-task transferability of performance occurring. In schema and knowledge triples, training on different tasks seems to weakly improve the performance of the evaluation task compared to finetuning on only that evaluation task. On tables, we see that this effect is much greater. One explanation for this is that there are a wide variety of table tasks included in the training mix, and the volume of auxiliary data is larger than the other data types.
On the other hand, we see that finetuning on different datatypes with the same task (i.e. summarization) also yields benefits to performance. On the summarization and question-answering (QA) experiments, we train on both tabular and knowledge graph data. We evaluate summarization with Dart and QA with WikiSQL. We see that even though we have only added one extra dataset in each case, we still see an improved result compared to the single-task model even though these datasets are have very different knowledge types. Thus, overall evidence supports that the performance of tasks within our training set are benefiting from each other, even across data types.
6 Discussion
We argue that SKG is an important capability for future language models. We have seen through our experiments on ChatGPT and the LLama2 family that there is significant room for improvement.
We found that we could produce a strong model by focused instruction-tuning on SKG tasks, however, we also observe that the performance difference between 7B to 34B StructLM models was not dramatic. This raises a concern about the state of SKG data: could we be approaching a performance ceiling? Combined with the fact that we were able to outperform UL2-20b, a much larger model, with our 7B model on 3 tasks, it seems that LLMs at various scales are struggling with SKG capabilities.
Indeed, grounding to structured knowledge directly in a model’s input represents a challenge in reasoning and input sensitivity. However, it has a wide range of potential benefits. To meaningfully improve SKG capability, we propose that future work may explore continued pretraining of open foundation models on more structured data formats. Similar to current attempts at code or math pretraining, it is possible that pretraining models on text interleaved with tables or other types of regular data formatting will help us move towards establishing SKG as a foundational model capability.
7 Conclusion
In this paper, we explore the current capabilities of open language models on structured knowledge grounding tasks. We show that LLMs are currently weak at SKG tasks currently. To address this gap, we construct an instruction-tuning dataset mixture of 1.1M examples and release models that outperform USKG on a majority ($14$ of $18$) tasks and achieve SOTA on $7$ of them. We also study the effects of various factors that influence the performance of our model on these task types. We hope that our work provides an updated understanding of what is achievable in the SKG domain, and can serve as a strong baseline for future improvements.
Limitations
The collection process used to construct the training data for StructLM tries to include a wide a variety of data types. As we have seen, there is evidence that this diversity is capable of affording transferable benefits to each dataset in the mixture. However, the tasks that we train and evaluate on are still academic datasets which have each been curated and designed for a specific purpose. While these are indeed diverse, the SKG domain relies on specific formatting and prompting conventions, which may result in our models having unnecessary specificity towards the conventions within our train set. To develop a clearer picture of how SKG performs as its own domain, we may require larger scale datasets with more heterogeneous formatting conventions. Further opportunities for training more robust SKG models may lie in increasing the diversity of structured data types in this way.
Additionally, while we have tried to evaluate our models to the best of our ability, many of the tasks of our held-out datasets measure accuracy through a heuristic matching step of a model’s output. In zero or few-shot settings, it is quite challenging to exactly control the generations of an autoregressive transformer to be adherent to a certain rule or grammar, and this has been a subject of study in other works (i.e. PICARD Scholak et al. (2021)). We note that because of this reality, poor results in zero or few-shot context may betray the existence of useful representations that the model has already learned. Without further prompting or finetuning efforts, it may be difficult to bring these capabilities to light. As such, another opportunity for improvement upon our methods may involve more flexible constrained methods of language model evaluation.
Ethics Statement
This paper does not raise any ethical concerns. The data and additional resources employed in this study are open-source and widely utilized in existing works.
Appendix A Appendix
Dataset Metric Split SOTA Score Best Performing Model
TabMWP Acc test 94.7 CREATOR (ChatGPT)(Qian et al., 2023)
ToTTo BLEU test 49.9 UniK2G (Li et al., 2024)
GrailQA EM val 77.1 TIARA + GAIN (T5-3B) (Gu et al., 2023)
SQL2Text Blec test 94.78 UnifiedSKG
MMQA F1 test 85.28 UnifiedSKG
Spider EM dev 80.5 RESDSQL (Li et al., 2023a)
KVRet All Micro test 67.88 UnifiedSKG
HybridQA Acc dev 68.4 S3HQA(Lee et al., 2023)
SParC EM test 68.2 CQR-SQL(Qi et al., 2022b)
CompWebQ Acc test 76.8 ChatKBQALuo et al. (2023b)
TabFact Acc test_small 93.0 Dater(Ye et al., 2023)
WikiTQ All Ex test 65.9 Dater(Ye et al., 2023)
WikiSQL All Ex test 93.0 SeaD+Execution-Guided Decoding(Xu et al., 2021)
FeTaQA BLEU test 39.0 TableLLama (Zhang et al., 2023)
Feverous Acc dev 85.6 FLAN UL2 20b(Tay et al., 2023b)
MultiWOZ Joint Acc test 60.6 TripPy + SaCLog(Dai et al., 2021)
Dart BLEU test 52.0 Control Prefixes (T5-large)(Clive et al., 2021)
Logic2Text Blec test 95.3 UnifiedSKG
MTOP EM test 87.5 FLAN UL2 20b(Tay et al., 2023b)
BIRD Acc dev 36.6 ChatGPT + COT(Li et al., 2023b)
CoSQL EM test 58.3 CQR-SQL(Xiao et al., 2022)
SQA Acc test 70.5 FLAN UL2 20b(Tay et al., 2023b)
Infotabs Acc dev 75.6 Infotabs paper(Gupta et al., 2020a)
WikiTableText BLEU test 33.7 UnifiedK2G(Li et al., 2024)
Finqa Acc private_test 71.1 APOLLO(Sun et al., 2022)
Table 4: Performance comparison of various models on different datasets, highlighting the SOTA scores and their sources. When selecting SoTA scores, we exclude methods that use large closed-source models such as GPT4 with agent-based methods. We also specify the data split that this score was achieved on.
Metric 0% 10% 20% 50% 57%
Held-In Datasets
TabMWP Acc 71.14 70.35 70.52 69.01 69.36
ToTTo BLEU 49.78 49.51 49.47 49.31 49.38
GrailQA EM 81.09 80.46 80.29 80.89 80.38
SQL2Text Blec 95.07 94.39 94.49 94.97 93.81
MMQA F1 84.26 84.31 84.11 83.40 85.15
Spider EM 72.92 71.57 73.40 72.73 72.44
KVRet All Micro 71.60 73.90 70.34 72.25 72.61
HybridQA Acc 59.23 59.09 59.09 59.03 59.17
SParC EM 63.09 62.34 63.26 64.59 61.93
CompWebQ Acc 80.61 79.15 78.76 78.73 78.34
TabFact Acc 83.41 81.09 81.42 80.92 80.77
WikiTQ All Ex 50.02 48.50 49.24 48.30 50.09
WikiSQL All Ex 87.33 86.45 86.73 86.68 88.67
FeTaQA BLEU 36.58 37.26 36.55 36.72 36.03
Feverous Acc 85.02 84.13 84.11 83.73 84.41
MultiWOZ Joint Acc 54.66 54.10 53.73 53.92 54.49
Dart BLEU 61.38 61.89 61.08 62.24 62.24
Logic2Text Blec 88.83 89.47 89.19 90.57 88.92
MTOP EM 82.44 81.71 81.19 80.92 81.21
Held-Out Datasets
BIRD Acc 21.30 22.30 22.30 23.00 22.30
CoSQL EM 51.24 49.95 50.84 50.74 49.75
SQA Acc 49.02 46.03 43.11 48.39 49.72
Infotabs Acc 38.00 56.26 57.87 62.35 62.46
WikiTableText BLEU 14.78 13.51 6.66 7.27 8.27
Finqa Acc 19.70 24.32 27.55 25.37 27.29
Table 5: Ablation results for the mixtures of general data in the training set. In total, we train 5 models, where the percentage represents the percent of the training data that is general. In the held out data, we see noticeable gains in generalization performance for FinQA and InfoTabs datasets. Notably, FinQA requires the generation of a python-executable math expression and InfoTabs requires an exact match to 3 previously unseen (boolean) options. WikiTableText performance seems to suffer, but is evaluated based the BLEU score with only one target sentence. As a result, we place more emphasis on the model’s 0-shot adaptation ability to new output specifications unseen in the training data.
Tasks Metric Code-LM LLaMA Math-LM
Held-In Datasets
TabMWP Acc 71.14 62.96 66.5
ToTTo BLEU 49.78 48.26 47.4
GrailQA EM 81.09 75.72 77.66
SQL2Text Blec 95.07 94.49 94.58
MMQA F1 84.26 83.96 82.13
Spider EM 72.92 65.96 71.95
KVRet All Micro 71.6 70.36 70.03
HybridQA Acc 59.23 59.26 57.04
SParC EM 63.09 56.94 60.35
CompWebQ Acc 80.61 77.31 76.6
TabFact Acc 83.41 80.46 79.47
WikiTQ All Ex 50.02 45.6 46.89
WikiSQL All Ex 87.33 83.93 85.49
FeTaQA BLEU 36.58 34.37 34.1
Feverous Acc 85.02 83.2 82.52
MultiWOZ Joint Acc 54.66 55.43 53.79
Dart BLEU 61.38 61.52 61.24
Logic2Text Blec 88.83 88.0 90.38
MTOP EM 82.44 77.18 75.56
Held-Out Datasets
BIRD Acc 21.3 15.9 18.8
CoSQL EM 51.24 42.8 48.76
SQA Acc 49.02 37.03 49.05
Infotabs Acc 38.0 4.44 32.54
WikiTableText BLEU 14.78 13.0 14.82
Finqa Acc 19.7 6.63 21.53
Table 6: The evaluation results comparing finetuning done on different base models. Code refers to Codellama-Instruct-7B. Math refers to LLemma-7b. LLaMA refers to Llama-2-7b.
[INST] <<SYS>> You are an AI assistant that specializes in analyzing and reasoning over structured information. You will be given a task, optionally with some structured knowledge input. Your answer must strictly adhere to the output format, if specified. <</SYS>> {instruction} {input} [/INST]
Figure 6: Prompt format for all SKG examples. This formatting convention follows LLama2 Touvron et al. (2023). The input contains the linearized structured data, together with any other context, question or statement. The instruction specifies the task.